OpenStudy (anonymous):

I need someone to explain how to solve this problem. What is the solution of the following system? 3x + 3y = 10 -9x - 9y = -30

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

multiply first equation by -3 what do you get ?

OpenStudy (anonymous):

Multiply all numbers by 3?

OpenStudy (anonymous):

@satellite73

ganeshie8 (ganeshie8):

yes multiplying an equation both sides by a number will not change the equation

OpenStudy (anonymous):

Ok, so how would I get the -3?

ganeshie8 (ganeshie8):

take first equation \[3x + 3y = 10\] multiply both sides by "-3", you get \[(-3)3x + (-3)3y = (-3)10\] simplying gives \[-9x-9y =-30\]

ganeshie8 (ganeshie8):

which is exactly same as the second equation

OpenStudy (anonymous):

Ahh, ok.

ganeshie8 (ganeshie8):

in light of that observation, what do you conclude ?

ganeshie8 (ganeshie8):

can we conclude that both the given equations represent the same line ?

OpenStudy (anonymous):

The solution to the system is -3. And yes we can seeing as though the first equation is equal to the second.

ganeshie8 (ganeshie8):

the solution is not -3

ganeshie8 (ganeshie8):

both the lines are same so there will be many many infinitely many solutions

OpenStudy (anonymous):

Oh ok ok. Thank you for helping me. I was at a loss.

ganeshie8 (ganeshie8):

yw!